OpenStudy (anonymous):

Solve the system using elimination. 5x + 4y = 12 3x – 3y = 18

OpenStudy (anonymous):

5x + 4y = 12 3x – 3y = 18 multiply top by 3 and bottom by 4

OpenStudy (anonymous):

(5x + 4y = 12)3......15x+12y=36 (3x – 3y = 18)4........12x-12y=72